Solve the system by substitution.
3x – y = –7
x + y = –9

Answer by nateobot(1)

3x – y = –7
x + y = –9
y = -9 -x


3x - (-9 - x) = -7
3x + 9 + x = -7
4x + 9 = -7
4x = -16
x = -4


-4 + y = -9
y = -5


Proof:


3(-4) - (-5) = -7
-12 + 5 = -7
-7 = -7


-4 + (-5) = -9
-9 = -9